The bumper cars crash into each other and stop.Explain why both bumper cars stop after the crash.

1
Expert's answer
2021-10-13T15:40:53-0400

They both stop (at frontal bump) because their mass is equal and, according to conservation of momentum, the total momentum must be 0. Two stopped cars have the total momentum of 0 (the sum of two equal masses times 0 velocity). Two cars of equal mass moving at the same speed in opposite directions also has zero (initial) momentum:


pi=mvmv=0,pf=m0+m0,pf=pi.p_i=mv-mv=0,\\ p_f=m·0+m·0,\\ p_f=p_i.